OpenStudy (anonymous):

what is sin(135) what is sin(135) @Mathematics

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

subtract 135 from 180 and that gives you the angle but in the first quadrant. So 180 - 135 = 45.. And sin(45) = (sqrt(2))/(2)

OpenStudy (anonymous):

tyler can you help me with sin (-13pi/4)